OpenStudy (anonymous):

A cruise ship can cover 19 nautical miles in 399 minutes. How many nautical miles will it travel in 231 minutes? A. 11 B. 15 C. 9 D. 13

OpenStudy (anonymous):

19/399 = x/231 cross multiply 399x = (231)(19) 399x = 4389 x = 4389/399 x = 11